OpenStudy (anonymous):

solve the following second degree equation using the quadratic formula. a+bx+cx^2=0 ; where c is not equal to zero.

OpenStudy (anonymous):

lol...

OpenStudy (saifoo.khan):

whts LOL for?

OpenStudy (anonymous):

x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}

OpenStudy (anonymous):

khuram try completing squares ! ; )

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] for ax^2 +bx +c so just switch your coefficients about and u r done.

OpenStudy (anonymous):

it is lol

OpenStudy (anonymous):

It would be different if asked to "derive" the formula itself, however...

OpenStudy (anonymous):

If it just says "solve" then u can assume the formula.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

the quesiont itself was a joke :P but can't change the coefficients :P

OpenStudy (anonymous):

a might be zero?:-)

OpenStudy (anonymous):

you are taking it wrong standard equation is ax^2 +bx +c = 0 and in your equation it is cx^2 +bx +a = 0

OpenStudy (anonymous):

yeah i know that standard equation is what you said! but this question is like cx^2 +bx +a = 0

OpenStudy (anonymous):

So swap the coefficients in the formula (and specify a not 0)

OpenStudy (anonymous):

so ..the x is now \[\frac{-b \pm \sqrt{b^2 - 4ac}}{2c}\]

OpenStudy (anonymous):

all estudier said was to swap a and c

OpenStudy (anonymous):

and maybe you were not able to solve this question ... and when estudier typed lol and me too and all..you felt something different( i don't want it to say ) and now you are kinda covering up (No offence )....I am telling you No Offence ..sorry if it hurts

OpenStudy (anonymous):

I didn't mean to cause any offence, please forgive me if i did..

OpenStudy (anonymous):

its okay man, i was just confirming my answer.